OpenStudy (anonymous):

For which 2 of the basic trig identities is one side of the equation defined at \(\theta=0\) and the other side isn't?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

anyone??

jimthompson5910 (jim_thompson5910):

for it to be an identity, it would have to be an equation that is defined for all allowed input values so to say that it's defined on one side, but not on the other, doesn't make sense because that would invalidate the fact it's an identity

jimthompson5910 (jim_thompson5910):

but if you're asking which trig functions are undefined at theta = 0, they are csc(theta) and cot(theta) because csc(x) = 1/sin(x) cot(x) = cos(x)/sin(x) and sin(x) = 0 when x = 0

OpenStudy (anonymous):

i dont know but the exact question from my book is: For exactly teo of the basic identities, one side of the equation is defined at \(\theta=0\) and the other side is not. Which two?

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

oh i see what it means, if you check out this page (which I'm assuming what your book looks like) http://fhs.fusdaz.org/download.axd?file=f908ec34-dd0e-44c2-b0aa-e39839832177&dnldType=Resource they list the basic identities of which you use to answer this question

jimthompson5910 (jim_thompson5910):

what I'm referring to is this

OpenStudy (anonymous):

yes i have that but i cant seem to figure it out

jimthompson5910 (jim_thompson5910):

so which one of these identities have one side defined at theta = 0, but theta = 0 is undefined on the other side?

OpenStudy (anonymous):

okay i think its \[sin\theta={1\over csc\theta}\]and\[tan\theta={1\over cot\theta}\]

OpenStudy (anonymous):

am i right?

jimthompson5910 (jim_thompson5910):

correct, sine is defined at theta = 0, but csc is not same with tan and cot

OpenStudy (anonymous):

okay one more question

OpenStudy (anonymous):

how do you do:\[{tan^2x\over sec~x+1}\]

OpenStudy (anonymous):

ir says to write the expression as an algebraic expression of a single trigonometric function....and it says the answer is \(sec~x+1\)

jimthompson5910 (jim_thompson5910):

from this page, http://www.sosmath.com/trig/Trig5/trig5/trig5.html 1 + tan^2x = sec^2x tan^2x = sec^2x - 1 tan^2x = (secx + 1)(secx - 1)

OpenStudy (anonymous):

then shouldnt the answer be secx-1 not +1?

jimthompson5910 (jim_thompson5910):

there must be a typo then

OpenStudy (anonymous):

in \[sin~x~tan^2x-sin~x=0\]how do you factor out sin x?

jimthompson5910 (jim_thompson5910):

like this sin(x)*tan^2(x) - sin(x) = 0 sin(x) [ tan^2(x) - 1 ] = 0

OpenStudy (anonymous):

im stupid..

jimthompson5910 (jim_thompson5910):

no you're not, just a silly mistake

OpenStudy (anonymous):

okay so how do you go on from there:\[sin~x=0~~~~;~~~~tan^2x=0\]it asks for what values of x sin x would equal 0...i dont get it

OpenStudy (anonymous):

tan^2(x)-1=0* sorry

jimthompson5910 (jim_thompson5910):

when is sin(x) = 0 true

OpenStudy (anonymous):

yea exactly...i dont get how you would find that

OpenStudy (anonymous):

wait is it 1 and -1?

jimthompson5910 (jim_thompson5910):

do you have the unit circle with you?

OpenStudy (anonymous):

i know it, yea

OpenStudy (anonymous):

sin x = 0 when x=+/-1

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

use this http://www.math.ucsd.edu/~jarmel/math4c/Unit_Circle_Angles.png

jimthompson5910 (jim_thompson5910):

locate the angles when the y coordinate is 0

OpenStudy (anonymous):

pi/2 and 3pi/2?

OpenStudy (anonymous):

yes??

jimthompson5910 (jim_thompson5910):

close, you're seeing the x coordinate as zero there though it's actually 0 and pi

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

and what about the other one?\[tan^2x-1=0~~\implies~~~tan~x=\pm~\underline{~~~~~~~~}?\]

OpenStudy (anonymous):

\[\pm1\]right?

jimthompson5910 (jim_thompson5910):

tan^2x-1 = 0 (tanx - 1)(tanx + 1) = 0 tanx - 1 = 0 or tanx + 1 = 0 tanx = 1 or tanx = -1 so yes, plus/minus 1

OpenStudy (anonymous):

okay and can you check y work for this one:\[2~sin^2x+3~sin~x=2\]\[2~sin^2x+3~sin~x-2=0\]\[(2~sin~x-1)(sin~x+2)=0\]\[2~sin~x-1=0~~~~~~~~~~~~~~sin~x+2=0\]\[sin~x={1\over2};{-2}\]yea?

jimthompson5910 (jim_thompson5910):

you got it, just solve each for x from here

OpenStudy (anonymous):

so that would be pi/6 and...whats the other one?

OpenStudy (anonymous):

isnt sin x=-2 a no solution since sin x can only be between -1 and 1 ??

jimthompson5910 (jim_thompson5910):

correct, there are no solutions to sin(x) = -2

jimthompson5910 (jim_thompson5910):

sin(x) = 1/2 has the solutions x = pi/6 and x = 5pi/6 the other solution is found by subtracting pi/6 from pi pi - pi/6 = 6pi/6 - pi/6 = 5pi/6

jimthompson5910 (jim_thompson5910):

this is assuming 0 < x < 2pi

OpenStudy (anonymous):

wouldnt it be 13pi/6?

jimthompson5910 (jim_thompson5910):

how are you getting that?

OpenStudy (anonymous):

on my sheet it says to add 2pi to the solution to get all the values of x

jimthompson5910 (jim_thompson5910):

oh gotcha

jimthompson5910 (jim_thompson5910):

well that just means add multiples of 2pi, so one is pi/6 + 2pi*k where k is any integer another is 5pi/6 + 2pi*k where k is any integer

OpenStudy (anonymous):

yea thats what it says

jimthompson5910 (jim_thompson5910):

you just leave it like that though

OpenStudy (anonymous):

okay and if you were to represent this expression involving only sines and cosines would this be it?\[{{(sec~y-tan~y)(sec~y+tan~y)}\over{sec~y}}~~~~=~~~~{cos~y}\]

OpenStudy (anonymous):

yes??

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

k

jimthompson5910 (jim_thompson5910):

sry got distracted, but yes that's correct

jimthompson5910 (jim_thompson5910):

the numerator turns into sec^2 - tan^2, which just 1

OpenStudy (anonymous):

lol no problem...cool :)

OpenStudy (anonymous):

hey brb i have a few more...thank you sooo much so far

OpenStudy (anonymous):

if tan x = (pi/2 - theta) = -5.32, find cot theta.

OpenStudy (anonymous):

and it gives the blanks:\[Since~~tan~x=({\pi\over2}-\theta)=~\underline{~~~~~~~~~~}~~then~~cot\theta=\underline{~~~~~~~~~}\]

jimthompson5910 (jim_thompson5910):

same page http://www.sosmath.com/trig/Trig5/trig5/trig5.html

jimthompson5910 (jim_thompson5910):

look where it says "Co-Function Identities"

OpenStudy (anonymous):

so is that\[tan~x=({\pi\over2}-\theta)=-5.32\]\[cot\theta=-{1\over5.32}\]yes?

jimthompson5910 (jim_thompson5910):

no, tan(pi/2 - x) = cot(x) according to the page

OpenStudy (anonymous):

ohhhh right sorry i didnt look at that ...so it would be cot theta=-5.32?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

take (sin x)^2 + (cox x)^2 = 1 and divide through by (cos x)^2, what do you get? take (sin x)^2 + (cox x)^2 = 1 and divide through by (sin x)^2, what do you get? i got (sec x)^2 and (csc x)^2

jimthompson5910 (jim_thompson5910):

(sin x)^2 + (cox x)^2 = 1 turns into tan^2 + 1 = 1/cos^2 after dividing everything by cos^2 which is really tan^2 + 1 = sec^2 (something we've already seen and used before in this post) ------------------------------------------------------- (sin x)^2 + (cox x)^2 = 1 turns into 1 + (cot x)^2 = 1/sin^2 after dividing everything by sin^2 which is really 1 + cot^2 = csc^2 these two identities are listed on that page I gave you

OpenStudy (anonymous):

so its what i said?

jimthompson5910 (jim_thompson5910):

yeah part of what you said is correct, but i think they wanted those identities

OpenStudy (anonymous):

im a little confused with what you did up there...i cant find that anywhere on the page you gave

jimthompson5910 (jim_thompson5910):

look for "Pythagorean Identities"

jimthompson5910 (jim_thompson5910):

second group on that page

OpenStudy (anonymous):

yea but none of those say what you said

jimthompson5910 (jim_thompson5910):

i just left off the variable

OpenStudy (anonymous):

okay can you explain it again a little simpler or something,idk i didnt get it the first time

jimthompson5910 (jim_thompson5910):

(sin x)^2 + (cos x)^2 = 1 (sin x)^2/(cos x)^2 + (cos x)^2/(cos x)^2 = 1/(cos x)^2 (tan x)^2 + 1 = (sec x )^2

jimthompson5910 (jim_thompson5910):

in that I'm dividing everything by (cos x)^2

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

yeah i just did it without the variables

OpenStudy (anonymous):

okay so its {(tan x)^2 + 1 = (sec x )^2}/(cos x)^2

jimthompson5910 (jim_thompson5910):

how are you getting that?

OpenStudy (anonymous):

when it says to divide by (cos x)^2

jimthompson5910 (jim_thompson5910):

yeah but when it turns into sec x, the cosine on the right side goes away

jimthompson5910 (jim_thompson5910):

so it's just (tan x)^2 + 1 = (sec x )^2

OpenStudy (anonymous):

oh so thats the answer?

jimthompson5910 (jim_thompson5910):

yeah for that first part

OpenStudy (anonymous):

and then the second one would be 1+ (cot x)^2 = (csc x)^2

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

okay yay! tan theta = 3 and cos theta > 0 since (tan theta)^2 +1 = (sec theta)^2 then 9 + 1 = (sec theta)^2, 10= (sec theta)^2, sqrt10= sec theta since cos theta > 0 then sec theta = -10 so cos theta = -1/10 yes?

OpenStudy (anonymous):

that was the last one. is that right?